OpenStudy (anonymous):

Please help! Simplify each rational expression. State any restrictions on the variables.

OpenStudy (anonymous):

\[\frac{ 12-x-x ^2}{ x^2-8x+15 }\]

OpenStudy (jdoe0001):

\(\large{ \cfrac{ 12-x-x ^2}{ x^2-8x+15 }\implies \cfrac{ -(x^2+x-12)}{ x^2-8x+15 } \\ \quad \\ --------------------\\ \begin{array}{cccllll} x^2&+x&-12\\ &\uparrow &\uparrow \\ &4-3&4\cdot -3 \end{array}\qquad \qquad \begin{array}{cccllll} x^2&-8x&+15\\ &\uparrow &\uparrow \\ &-3-5&-3\cdot -5 \end{array}}\)

OpenStudy (anonymous):

Could you explain that to me? I don't get it.

OpenStudy (jdoe0001):

have you done quadratic factoring yet?

OpenStudy (anonymous):

Just a little bit @jdoe0001